# An X-Ray High-Frequency QPO in NGC 1365

### v1

@@
An X-Ray High-Frequency QPO in NGC 1365
Yongkang Yan 11    Peng Zhang E-mail:zhangpeng@ctgu.edu.cn11 2 2    Qingzhong Liu 33    Zhi Chang 44    Gaochao Liu 11 2 2    Jingzhi Yan 33    Xiangyun Zeng 11 2 2
@@
This study presents the detection of a high-frequency Quasi-Periodic Oscillation (QPO) in the Seyfert galaxy NGC 1365, based on observational data obtained by the XMM-Newton in January 2004. Utilizing the Weighted Wavelet Z-transform (WWZ) and Lomb-Scargle Periodogram (LSP) methods, a QPO signal was identified at a frequency of $2.19\times 10^{-4}\ {\rm Hz}$ (4566 s), with a confidence level of 3.6 $\sigma$. The signal was notably absent in the lower 0.2 – 1.0 keV energy band, with the primary contribution emerging from the 2.0 – 10.0 keV band, where the confidence level reached 3.9 $\sigma$. Spectral analysis shows that there are multiple absorption and emission lines in the high-energy band (¿ 6 keV). The correlation between the QPO frequency ($f_{\rm QPO}$) and the mass of NGC 1365 central black hole ($M_{\rm BH}$) aligns with the established logarithmic trend observed across black holes, indicating the QPO is of high frequency. This discovery provides new clues for studying the generation mechanism of QPO in Seyfert galaxies, which helps us understand the accretion process around supermassive black holes and the characteristics of strong gravitational fields in active galactic nuclei.
@@ 1 Introduction
Active Galactic Nuclei (AGN) are a class of extragalactic objects characterized by intense energetic processes occurring at their cores, primarily driven by the accretion of matter onto Supermassive Black Holes (SMBHs). The presence of anisotropic structures within the galactic nuclei bestows a variety of observational characteristics across the electromagnetic spectrum (Swain et al., 2023), with the X-ray spectrum being notably shaped by the re-radiation resulting from the interaction of X-ray photons with the surrounding gas. This X-ray radiation is believed to originate from regions within the accretion disk or closer to the black hole(Ghisellini et al., 1994). A defining feature of AGN X-ray emission is its temporal variability, which can be attributed to changes in the accretion processes and the shielding effects of the environment (Lachowicz & Czerny, 2005). Quasi-periodic oscillations are a special phenomenon of some X-ray and $\gamma$-ray emission sources, which are quite common in neutron stars and black hole binaries in the Milky Way and nearby galaxies (Gierlinski et al., 2008; Pan et al., 2016). However, this phenomenon is very rare in AGN (Zhang et al., 2020), and with many early detections have been disfavored due to inadequacies in modeling the underlying broad-band noise (Alston et al., 2015).
Based on the time variability and spectral changes of radiation, an outburst cycle can be divided into five stages: quiescent state, low hard state, hard intermediate state, soft intermediate state, and high soft state (Muñoz-Darias et al., 2011). In X-ray outbursts, periodicity reappears, and its frequency is not constant. For example, in one observation of RE J1034+396, there was a change in the QPO frequency before and after (Gierlinski et al., 2008). According to the frequency size, QPO can be roughly divided into low-frequency QPO (LFQPO) and high-frequency QPO (HFQPO) (Ma et al., 2021). For black hole X-ray binary, frequency below 30 Hz is considered LFQPO (Remillard & McClintock, 2006). However, for AGN, due to their large mass, the QPO frequencies that appear are generally low, mostly at the $10^{-4}$ Hz level (Kluzniak & Abramowicz, 2002; Zhou et al., 2010; Pan et al., 2016; Zhang et al., 2018). In addition, there is a close relationship between the mass of the central black hole of AGN and the QPO frequency. Zhou et al. (2015) believes that if the $f_{\rm QPO}$ of AGN has a certain relationship with $M_{\rm BH}$, it can be considered as HFQPO. The specific relationship will be explained in the discussion section.
In a wide range of studies, QPOs often exhibit transient characteristics (Gierlinski et al., 2008; Pan et al., 2016), which is a short-lived phenomenon. There are various models to explain its origin, but a unified theory has not yet been formed (Zhang et al., 2020). The explanations for the origin of different types of QPOs also vary greatly. For LFQPO, the most commonly applicable model is the relativistic precession model (Lense & Thirring, 1918), which was first calculated by Lense and Thirring in 1918 under the conditions of ignoring viscosity and weak field approximation (since then, this precession has been called LT precession). At present, the more complete relativistic precession model is improved by Ingram et al. (2009). The origin of HFQPO may be related to the orbital motion of the inner accretion flow of the accretion disk. The frequency of HFQPO is relatively stable and does not change much with the change of luminosity. The main explanatory models are the resonance model (Abramowicz & Kluźniak, 2001), the accretion-jet instability in the accretion disk (Tagger & Pellat, 1999), and disk oscillations in different modes (g, c, p mode) (Wagoner, 1999).
Regardless of these rapidly changing mechanisms, they occur very close to black holes, and black holes have the most direct impact on them. Therefore, by studying the QPO properties of AGN, we can better understand the accretion process from a constant black hole to a supermassive black hole and the accretion disk theory around it (Zhang et al., 2020). This also provides a window for us to study the strong gravitational field around the black hole.
@@
NGC 1365 is a large barred spiral galaxy in the Fornax galaxy cluster (Sandage & Tammann, 1983), with a distance of about 18.6 $\pm$ 0.6 Mpc (Madore et al., 1999; Silbermann et al., 1999; Springob et al., 2009; Polshaw et al., 2015; Jang et al., 2018). According to the latest calculations, its redshift value is z = 0.0051 (Koss et al., 2022). This spiral galaxy has been intensely investigated (a review of early work can be found in Lindblad (1999)). The galaxy contains an active core, and its spectral emission lines have both broad and narrow lines. The mass of the central black hole is in the range of (5 – 10) ×$\times 10^{6}\ M_{\odot}$ Maiolino & Rieke (1995), and the mass of the supermassive black hole can be derived from the emission lines in the broad-line region, in particular the broad components of hydrogen recombination lines, which can be calculated using the formula obtained in the near-infrared regionKim et al. (2010). NGC 1365 is also a typical Seyfert galaxy, and according to the Seyfert classification, it belongs to type 1.8 (Maiolino & Rieke, 1995). However, some scholars have classified it as type 1, 1.5 and 2 (e.g., Veron et al. (1980); Turner et al. (1993); Risaliti et al. (2007); Thomas et al. (2017)). This is due to the different classifications caused by the change in X-ray flux and the absorption body density along the line of sight (Hernández-García et al., 2015). The X-rays from the core include the hard continuous radiation of the active nucleus itself, the Fe-k line radiation of the rotating disk, and the thermal radiation of the surrounding stellar burst activity.
@@
In this work, we employ the Weighted Wavelet Z-transform (WWZ) and Lomb-Scargle Periodogram (LSP) to detect and characterize a QPO signal in the X-ray observations of NGC 1365. The study’s findings provide insights into the ongoing discussion regarding the transient nature of Quasi-Periodic Oscillations (QPOs) and their potential correlation with the properties of the central black hole. The structure of this paper is as follows: Section 2 details the observations, data reduction, and analysis methods, while Section 3 summarizes the results and discusses their implications for the study of AGN and the broader astrophysics community.
@@ 2.1 Observations and Data Reduction
Figure 1: Upper panel shows the light curve of the EPIC camera of the XMM-Newton ObsID 0205590301 in the range of 0.2 – 10.0 keV. Lower-left panel shows a two-dimensional plot of dynamic WWZ power in time-frequency space. Lower-right panel displays average WWZ power (black) and LSP value (red) curves. The blue lines represent the confidence level of 3.0, 3.5 and 4.0 $\sigma$, respectively. The orange dotted line represents 4.0 $\sigma$ confidence level calculated using the method proposed by Baluev (2008).
Figure 2: The pulse shape of the light curve is folded using period of 4566 s (two cycles are shown).
Figure 3: Part A shows WWZ power for ObsID 0205590301 at 0.2 – 1.0 keV, 1.0 – 2.0 keV, 2.0 – 6.0 keV and 6.0 – 10.0 keV. Part B, C, D and E represent the results of wavelet analysis at different energy bands, with the frequency range between two black dashed lines in Part A.
XMM-Newton is an X-ray detection satellite launched by the European Space Agency on December 10, 1999. It carries two sets of X-ray detectors, including three European Photon Imaging Cameras (EPIC; PN, MOS1, MOS2; Strüder et al. (2001); Turner et al. (2001)) and two Reflection Grating Spectrometers (2RGS;den Herder et al. (2001)). Within the span of 2003 and 2004 alone, XMM-Newton conducted 6 observations for NGC 1365. After a considerable interval, another observation was recently executed on February 21, 2024. In total, there have been 14 observations, with 10 of them being long-duration observations ($>$ 30 ks). In this paper, the observation data we used is from January 17, 2004, with an exposure time of 59.7 ks, and the observation number is ObsID 0205590301. We preprocess the observation data using the Scientific Analysis Software (SAS, version 21.0.0) provided by the XMM-Newton Science Operations Center, and extract the required scientific data using the evselect tool according to the standard procedure.
@@
Traditional tools for studying time variability include Power Spectral Density (PSD) analysis, Short-Time Fourier Transform (STFT) of light curves and LSP. PSD and LSP are best suited for searching for strictly periodic signals, not designed for analyzing signals with time-varying frequencies. STFT can handle variable frequencies, but it largely depends on the choice of window function, and if chosen improperly, it usually produces significant side effects (Zhao et al., 2020). WWZ has a robust mapping in the frequency-time domain, especially when detecting the time evolution of parameters (such as period, amplitude, phase) describing periodic and quasi-periodic signals, WWZ shows great advantages as a periodic analysis method (Zhang et al., 2020). Examples of WWZ analysis of the evolution characteristics of QPOs in various celestial systems include Bravo et al. (2014), Ait Benkhali et al. (2020), Urquhart et al. (2022).
In this paper, a version of the WWZ modified with the Morlet mother function is utilized. (Foster, 1996). A WWZ analysis is conducted on the obtained light curve, resulting in three sets of two-dimensional color maps that depict the power spectrum over time. By observing these two-dimensional maps, we can determine whether the light curve has a period and the location of the period. At the same time, LSP method is additionally employed for further analysis and to perform comparative checks. The LSP method is different from the principle method used by WWZ. Its advantage is that it can be used to analyze the periodicity in irregular time series, by fitting the sine waves in the entire data series using the $\chi^{2}$ statistic. This method reduces the impact of irregular sampling and points out any periods or quasi-periods that may exist in the data, and it can also calculate their significance.
@@
In addition, we also used the method proposed by (Baluev, 2008) to calculate the $4.0\ \sigma$ confidence level. Represented by the orange dotted line in Figure 1, you can see that the result is significantly more than 4.0 $\sigma$. However, this method does not take into account the impact of red noise, and for observations with a high impact of red noise, the results calculated by this method are somewhat controversial. Therefore, we can conclude that the QPO discovered in this study has a high level of confidence.
@@
The entire time span of the EPIC light curve was folded with a period of 4566 s, and the result is shown in Figure 2. The error of each point is calculated based on the standard deviation ($68.3\ \%$) of the mean values of each phase interval. For clarity, we plotted two periods. The average count rate is about $\sim$ 3.07 counts/s. From the figure, it can be seen that the amplitude of the X-ray flux changes significantly with the phase.
@@
In order to more accurately analyze which energy band the QPO originates from, we divided 0.2 – 10 keV into four parts: 0.2 – 1.0 keV, 1.0 – 2.0 keV, 2.0 – 6.0 keV and 6.0 – 10.0 keV. For different energy bands, we obtained the light curves using the same parameters and data processing methods. Their WWZ results(Figure 3) show that there was no QPO signal detected in 0.2 – 1.0 keV energy band, only a hint of presence at 1.0 – 2.0 keV, and the main contribution of the QPO signal comes from the 2.0 – 10.0 keV range. Similarly, to more clearly describe this difference, we calculated the confidence level for the 2.0 – 10.0 keV energy band. Through $10^{6}$ artificial light curves, we obtained a confidence level of 3.9 $\sigma$.
2.2 Time-averaged Spectra Analysis

### v2

# An X-ray high-frequency quasi-periodic oscillation in NGC 1365

@@
An X-ray high-frequency quasi-periodic oscillation in NGC 1365
Y. K. Yan 11    P. Zhang E-mail:zhangpeng@ctgu.edu.cn11 2 2    Q. Z. Liu 33    Z. Chang 44    G. C. Liu E-mail:gcliu@ctgu.edu.cn11 2 2    J. Z. Yan 33    X. Y. Zeng 11 2 2
@@
This study presents the detection of a high-frequency quasi-periodic oscillation (QPO) in the Seyfert galaxy NGC 1365 based on observational data obtained by XMM-Newton in January 2004. Utilizing the weighted wavelet Z-transform (WWZ) and Lomb-Scargle periodogram (LSP) methods, a QPO signal is identified at a frequency of $2.19\times 10^{-4}\ {\rm Hz}$ (4566 s), with a confidence level of 3.6 $\sigma$. The signal is notably absent in the lower 0.2 – 1.0 keV energy band, with the primary contribution emerging from the 2.0 – 10.0 keV band, where the confidence level reaches 3.9 $\sigma$. Spectral analysis shows that there are multiple absorption and emission lines in the high-energy band (¿ 6 keV). The correlation between the QPO frequency ($f_{\rm QPO}$) and the mass of the central black hole ($M_{\rm BH}$) of NGC 1365 aligns with the established logarithmic trend observed across black holes, indicating the QPO is of high frequency. This discovery provides new clues for studying the generation mechanism of QPOs in Seyfert galaxies, which helps us understand the accretion process around supermassive black holes and the characteristics of strong gravitational fields in active galactic nuclei.
@@ 1 Introduction
Active galactic nuclei (AGNs) are a class of extragalactic objects characterized by intense energetic processes occurring at their cores, which is primarily driven by the accretion of matter onto a supermassive black hole (BH). The presence of anisotropic structures within the galactic nuclei bestows a variety of observational characteristics across the electromagnetic spectrum (Swain et al., 2023), with the X-ray spectrum notably shaped by the reradiation resulting from interactions between X-ray photons and the surrounding gas. This X-ray radiation is believed to originate from regions within the accretion disk or closer to the black hole (Ghisellini et al., 1994). A defining feature of AGN X-ray emission is its temporal variability, which can be attributed to changes in the accretion processes and the shielding effects of the environment (Lachowicz & Czerny, 2005). Quasi-periodic oscillations (QPOs) are a special phenomenon of some X-ray and $\gamma$-ray emission sources, and are quite common in neutron stars and BH binaries in the Milky Way and nearby galaxies (Gierliński et al., 2008; Pan et al., 2016). However, this phenomenon is very rare in AGNs (Zhang et al., 2020), and many early detections have been disfavored due to inadequacies in modeling the underlying broadband noise (Alston et al., 2015).
An important observational property of stellar-mass BH X-ray binaries, potentially a probe of the inflow structure just outside the horizon, is the presence of QPOs (Remillard & McClintock, 2006). These phenomena exhibit different peaks in their X-ray power density spectrum. QPOs are usually divided into high frequency (HF; $\sim$10 – $10^{3}$ Hz) and low frequency (LF; $\sim$ 10-2 – 10 Hz). The latter are further classified into subtypes based on their coherence and the intensity of different frequency bands (Type A, B, and C: Casella et al. 2004). Coherence is represented by the quality factor Q $=\nu/\Delta\nu$, where $\nu$ is the centroid frequency and $\Delta\nu$ is the full width at half maximum near the centroid frequency. The most common QPO subtype (and also the subtype with the highest Q-factor) is type C; its frequency usually ranges from a few mHz to 10 Hz, but sometimes frequencies as high as 30 Hz can be detected (Revnivtsev et al., 2000). The explanation for the origin of type C QPO is still controversial; alternative models include internal disk instability and Lense Thirring (LT) precession of jets or internal heat flux or corona (Stella et al., 1999; Ingram et al., 2009; Ingram & Done, 2011).
These QPO phenomena exhibit similarities across compact objects of different masses and types, such as microquasars, AGNs, or intermediate-mass BHs. The correlation between HF QPOs and LF QPOs observed in neutron stars may also exist in weak galactic BH QPO data (see Motta et al. 2022). Although multiple models have attempted to explain the origin of these QPOs, no unified theory has yet been established (Zhang et al., 2020). Explanations for the origin of different types of QPOs vary significantly. For LF QPOs, Stella et al. (1999) proposed the so-called relativistic precession model, which links QPOs to the radial (periastron) and vertical (LT) precessions of inhomogeneities in the accretion disk, approximated as test particles. The origin of HF QPOs, on the other hand, may be related to the orbital motion of the accretion flow within the disk, with relatively stable frequencies that do not significantly change with luminosity. The main explanatory models include the resonance model (Abramowicz & Kluźniak, 2001), accretion-ejection instability within the disk (Tagger & Pellat, 1999), and disk oscillations in different modes (g, c, and p mode) (Wagoner, 1999).
Furthermore, the relationship between BH mass (M) and QPO frequency is of significant importance in the present study, as is its potential connection with the relativistic 1/M scaling of orbital frequencies. The expected 1/M scaling is generally valid for most orbital models, with recent key findings supporting this evidence (Šrámková et al., 2021). Additionally, extensive studies have shown that QPOs typically exhibit transient characteristics, indicating a short-lived phenomenon (Gierliński et al., 2008; Pan et al., 2016). Regardless of these rapidly changing mechanisms, they occur very close to the BH, which exerts the most direct influence on them. By studying the characteristics of QPOs, we can gain a better understanding of the accretion processes from stellar-mass to supermassive BHs and the accretion disk theory surrounding them (Zhang et al., 2020). This also provides a critical way to study the strong gravitational fields around BHs.
@@
NGC 1365 is a large barred spiral galaxy in the Fornax galaxy cluster (Sandage & Tammann, 1983), with a distance of about 18.6 $\pm$ 0.6 Mpc (Madore et al., 1999; Silbermann et al., 1999; Springob et al., 2009; Polshaw et al., 2015; Jang et al., 2018). According to the latest calculations, its redshift value is z = 0.0051 (Koss et al., 2022). This spiral galaxy has been extensively investigated (a review of early work can be found in Lindblad 1999). The galaxy contains an active core, and its spectral emission lines have both broad and narrow lines. The mass of the central BH is in the range of (5 – 10) ×$\times 10^{6}\ M_{\odot}$ (Fazeli et al., 2019), and the mass of the supermassive BH can be derived from the emission lines in the broad-line region, in particular the broad components of hydrogen recombination lines, which can be calculated using the formula obtained in the near-infrared (NIR) region (Kim et al., 2010). NGC 1365 is also a typical Seyfert galaxy, and according to the Seyfert classification, it belongs to type 1.8 (Maiolino & Rieke, 1995). However, some scholars have classified it as type 1, 1.5, and 2 (e.g., Veron et al. 1980; Turner et al. 1993; Risaliti et al. 2007; Thomas et al. 2017). This is due to the different classifications caused by the change in X-ray flux and the absorption body density along the line of sight (Hernández-García et al., 2015). The X-rays from the core include the hard continuous radiation of the active nucleus itself, the Fe-k line radiation of the rotating disk, and the thermal radiation of the surrounding stellar burst activity.
@@
In this work, we employ the weighted wavelet Z-transform (WWZ) and the Lomb-Scargle periodogram (LSP) to detect and characterize a QPO signal in the X-ray observations of NGC 1365. Our findings provide insights into the ongoing discussion regarding the transient nature of QPOs and their potential correlation with the properties of the central BH. The structure of this paper is as follows: Section 2 details our observations, data reduction, and analysis methods, while in Section 3 we summarize our results and discuss their implications for the study of AGNs and for the broader astrophysics community.
@@ 2.1 Observations and data reduction
Figure 1: Light curve and WWZ power of EPIC camera of the XMM-Newton ObsID 0205590301. The upper panel shows the light curve in the range of 0.2 – 10.0 keV. The upper panel shows the light curve of the EPIC camera of the ObsID 0205590301 in the range of 0.2 – 10.0 keV. The lower-left panel displays a two-dimensional plot of dynamic WWZ power in time–frequency space. The lower-right panel displays average WWZ power (black) and LSP value (red) curves. The blue lines represent the confidence levels of 3.0, 3.5, and 4.0 $\sigma$, respectively. The orange dotted line represents the 4.0 $\sigma$ confidence level calculated using the method proposed by Baluev (2008).
Figure 2: Pulse shape of the light curve, folded using a period of 4566 s, with two cycles shown.
Figure 3: WWZ power for ObsID 0205590301 at different energy band. Panel A shows average WWZ power for ObsID 0205590301 at 0.2 – 1.0 keV, 1.0 – 2.0 keV, 2.0 – 6.0 keV, and 6.0 – 10.0 keV. Panels B, C, D, and E represent the results of wavelet analysis at different energy bands, with the frequency range between two black dashed lines in Panel A.
XMM-Newton is an X-ray detection satellite launched by the European Space Agency on December 10, 1999. It carries two sets of X-ray detectors, including three European Photon Imaging Cameras (EPIC: PN, MOS1, and MOS2; Strüder et al. 2001; Turner et al. 2001) and two Reflection Grating Spectrometers (2RGS; den Herder et al. 2001). Within the span of 2003 and 2004 alone, XMM-Newton conducted six observations of NGC 1365. After a considerable interval, another observation was recently executed on February 21, 2024. In total, there have been 14 observations, with 10 of them being long-duration observations ($>$ 30 ks). In this paper, the observation data we used are from January 17, 2004, with an exposure time of 59.7 ks, and the observation number is ObsID 0205590301. We preprocess the observation data using the Scientific Analysis Software (SAS, version 21.0.0) provided by the XMM-Newton Science Operations Center, and extract the required scientific data using the evselect tool according to the standard procedure.
@@
Traditional tools for studying time variability include power spectral density (PSD) analysis, LSP, and short-time Fourier transform (STFT) of light curves. PSD and LSP are best suited to searching for strictly periodic signals, and are not designed for analyzing signals with time-varying frequencies. STFT can handle variable frequencies, but it largely depends on the choice of window function, and if chosen improperly, it usually produces significant side effects (Zhao et al., 2021). WWZ has a robust mapping in the frequency-time domain, especially when detecting the time evolution of parameters (such as period, amplitude, and phase) describing periodic and quasi-periodic signals, and shows great advantages as a periodic analysis method (Zhang et al., 2020). Examples of WWZ analysis of the evolution characteristics of QPOs in various celestial systems include Bravo et al. (2014), Ait Benkhali et al. (2020), and Urquhart et al. (2022).
In the present study, we used a version of the WWZ modified with the Morlet mother function (Foster, 1996). We conducted a WWZ analysis on the obtained light curve, obtaining three sets of two-dimensional color maps that depict the power spectrum over time. By observing these two-dimensional maps, we can determine whether or not the light curve has a period, and the location of the period. At the same time, we additionally employed the LSP method for further analysis and to perform comparative checks. The LSP method differs from the principle method used by WWZ. Its advantage is that it can be used to analyze the periodicity in irregular time series by fitting the sine waves in the entire data series using the $\chi^{2}$ statistic. This method reduces the impact of irregular sampling and identifies any periods or quasi-periods that may exist in the data, and it can also calculate their significance.
@@
In addition, we also used the method proposed by Baluev (2008) to calculate the $4.0\ \sigma$ confidence level. Represented by the orange dotted line in Figure 1, you can see that the result is significantly more than 4.0 $\sigma$. However, this method does not take into account the influence of red noise, and the results calculated using this method are somewhat controversial for observations that are greatly affected by red noise. Therefore, we can conclude that the QPO discovered in this study has a high level of confidence. In addition to a very obvious peak at a frequency of $2.19\times 10^{-4}\ {\rm Hz}$ in Figure 1, there seems to be signals at frequencies of $2.66\times 10^{-4}\ {\rm Hz}$ and $5.46\times 10^{-4}\ {\rm Hz}$ in the first $\sim$20 ksec. We pointed out in Zhang et al. (2023) that QPO signals do not persist indefinitely, but rather have a lifespan. This signal may occur at the end of a QPO signal before observation, or it may be a false signal that we cannot make judgments about based on existing information. We can therefore only calculate the confidence level of its signal based on limited information. In this case, we rule out the possibility that these are real signals.
@@
The entire time span of the EPIC light curve was folded with a period of 4566 s, and the result is shown in Figure 2. The error of each point is calculated based on the standard deviation ($68.3\ \%$) of the mean values of each phase interval. For clarity, we plotted two periods. The average count rate is $\sim$ 3.07 counts/s. From the figure, it can be seen that the amplitude of the X-ray flux changes significantly with the phase.
@@
In order to more accurately distinguish the energy band from which the QPO originates, we divided 0.2 – 10 keV into four parts: 0.2 – 1.0 keV, 1.0 – 2.0 keV, 2.0 – 6.0 keV, and 6.0 – 10.0 keV. For different energy bands, we obtained the light curves using the same parameters and data processing methods. Their WWZ results (Figure 3) show that there was no QPO signal detected in the 0.2 – 1.0 keV energy band, and only the hint of a signal at 1.0 – 2.0 keV, and the main contribution of the QPO signal comes from the 2.0 – 10.0 keV range. Similarly, to more clearly describe this difference, we calculated the confidence level for the 2.0 – 10.0 keV energy band. Through $10^{6}$ artificial light curves, we obtained a confidence level of 3.9 $\sigma$.
2.2 Time-averaged spectral analysis